use dimensional analysis and show tht C.G.S unit of density is 1000 times large as its units.....................................this is the question..........how to do this?? i tried a lot bt failed to get it ........wanted the answer today itself as tomorrow is my test...

Grade:Upto college level

1 Answers

AskiitianExpert Shine
14 years ago

Hi

S.I unit of density is Kg/m3.

1Kg= 1000gram  : gram is cgs unit

1m = 100 cm  : cm is cgs unit

therefore cgs unit is g/cm= 1o-3 kg/ 10-6 m3 = 1/1000 kg/m3 , hence proved.